If x+1;x-1 and x are the first three terms of a geometric sequence , find x and first three terms of the sequence .
1
Expert's answer
2019-08-22T11:50:54-0400

Since the first three terms of a geometric sequence are "x+1" ,"x-1" , "x" , the middle term must be the geometric mean of the other two. Hence

"(x-1)^2 =(x+1)\cdot x"

"x^2-2x+1=x^2 +x"

"-2x-x=-1"

"-3x=-1"

"x= \frac {1}{3}"

The first term is "x+1=\frac{1}{3}+1= 1\frac{1}{3} =\frac{4}{3}".

The second term is "x-1=\frac{1}{3}-1= -\frac{2}{3}"

The third term is "x=\frac{1}{3}"

Answer: the first three terms of a geometric sequence are "\frac{4}{3}, -\frac{2}{3}, \frac{1}{3}"